### v2

# GWQ: Gradient-Aware Weight Quantization for Large Language Models

Abstract
Large language models (LLMs) show impressive performance in solving complex language tasks. However, its large number of parameters present significant challenges for the deployment and application of the model on edge devices. Compressing large language models to low bits can enable them to run on resource-constrained devices, often leading to performance degradation. To address this problem, we propose gradient-aware weight quantization (GWQ), the first quantization approach for low-bit weight quantization that leverages gradients to localize outliers, requiring only a minimal amount of calibration data for outlier detection. GWQ retains the weights corresponding to the top 1% outliers preferentially at FP16 precision, while the remaining non-outlier weights are stored in a low-bit format. GWQ found experimentally that utilizing the sensitive weights in the gradient localization model is more scientific compared to utilizing the sensitive weights in the Hessian matrix localization model. Compared to current quantization methods, GWQ can be applied to multiple language models and achieves lower PPL on the WikiText2 and C4 dataset. In the zero-shot task, GWQ quantized models have higher accuracy compared to other quantization methods. GWQ is also suitable for multimodal model quantization, and the quantized Qwen-VL family model is more accurate than other methods. Zero-shot target detection task dataset RefCOCO outperforms the current stat-of-the-arts method SPQR. GWQ achieves 1.2$\times$ inference speedup in comparison to the original model, and effectively reduces the inference memory.
GWQ: Gradient-Aware Weight Quantization for Large Language Models
Yihua Shao1,2, Siyu Liang3, Zijian Ling3, Minxi Yan3, Haiyang Liu4, Siyu Chen2, Ziyang Yan5, Chenyu Zhang5, Haotong Qin6*, Michele Magno6, Yang Yang2, Zhen Lei2, Yan Wang3, Jingcai Guo7, Ling Shao8, Hao Tang1* 1PKU  2CASIA  3THU  4USTB  5UNITN  6ETHz  7PolyU  8UCAS thanks: The first two authors contributed equally. correspondence authors: qinhaotong@gmail.com, haotang@pku.edu.cn
1 Introduction
Large language models (LLMs) (Touvron et al., 2023a; Achiam et al., 2023; Almazrouei et al., 2023; Touvron et al., 2023b) based on Transformer (Vaswani, 2017) have demonstrated their outstanding capabilities in scenarios such as complex question and answer, language modeling (Brown, 2020), and so on. It can often generate more accurate answers with different inputs. Its ability to handle complex linguistic tasks is due to its huge pre-trained corpus (Kaplan et al., 2020) and an astronomical number of parameters (Chowdhery et al., 2023). Its outstanding performance is rapidly generating huge applications in people’s lives. However, the huge memory consumption generated by its pre-training and inference phases leads to great difficulties in its application and deployment (Li et al., 2024a). Therefore, when deploying large language models, there are often huge GPU clusters to support the inference of the models (Xu et al., 2024).
In order to make large language models adaptable to resource-constrained edge devices (Li et al., 2024b; Huang et al., 2024), model compression (Ma et al., 2023; Gu et al., 2023) has become a common means to reduce the computational memory of models. Among them, post-training large language model quantization (Dettmers and Zettlemoyer, 2023) is a commonly used method to reduce the model inference overhead. Most current quantization methods compress the model to 3 or 4 bits (Dettmers et al., 2023; Frantar et al., 2022a), while trying to ensure that the model has as little loss of accuracy and performance as possible. Post-training quantization of pre-trained models reduces model inference overhead by allowing the number of parameter bits to decrease. This approach produces a loss of accuracy that affects the performance of the model in use. Therefore, to reduce the size of the model while ensuring its performance, it is necessary to design effective quantization algorithms (Lin et al., 2024). With effective quantization algorithms, researchers can make the performance of low-bit models infinitely close to that of 16-bit models.
OBQ (Frantar and Alistarh, 2022) indicates that when the pre-trained model has fully converged, the model should exhibit zero gradients. However, our experiments reveal that the LLMs still generate gradients in response to different text inputs, indicating that the model has not completely converged and remains sensitive to the input. Inspired by this phenomenon, we propose a gradient-aware post-training weight-only quantization method called GWQ. GWQ is the first post-training quantization approach to utilize gradients to locate outliers in pre-trained models. It introduces two key innovations. First, GWQ selects 1% of the weights as outliers by looking at the magnitude of the gradient for the response to the calibration set, preserving these outliers in FP16 precision while quantizing the remaining 99% of the weights to 4 bits or 3 bits.
In summary, the primary contributions of this paper are as follows:
  • 1)
    GWQ discovered that employing the first-order gradient to search for and locate sensitive weights is more rational than using the Hessian matrix for localization, and the model loss after quantization is also lower. Therefore, using a first-order gradient to locate outliers is more reasonable.
  • 2)
    GWQ is the first accurate first-order gradient-aware post-training weight quantization method for pre-trained LLMs, requiring only a minimal quantity of calibration data to identify outliers efficiently.
  • 3)
    GWQ outshines the current state-of-the-arts method SPQR on the wikitext and C4 datasets. The majority of zero-shot detection tasks in the multimodal dataset RefCOCO have reached the state-of-the-arts. Meanwhile, the quantified model has achieved a 1.2$\times$ acceleration compared to the original model and utilized less memory during inference.
2 Related Work
Post-training Quantization. Post-training quantization (PTQ) applies mainly to visual models (Gholami et al., 2022). For example, AdaRound (Nagel et al., 2020), AdaQuant (Hubara et al., 2020), BitSplit (Wang et al., 2024b) can be used for models with a small number of parameters. When models have a large number of parameters, they cause a significant loss of accuracy, and the reduction in memory overhead of the quantized model compared to the original model is not particularly significant. Most current quantization methods for models are improved by Optimal Brain Quantization (OBQ) (Frantar and Alistarh, 2022). OBQ is improved by the Optimal Brain Surgeon (OBS) (Hassibi et al., 1993; Singh and Alistarh, 2020; Frantar et al., 2021), both OBS and OBQ default to a model whose response to input is 0 at the end of training, but the OBQ method applies the idea of weight pruning to model quantization by quantifying the weights that have the least impact on the network and compensating for them with the remaining weights. Some current PTQ methods perform outlier search by reinforcement learning (Wang et al., 2024b) to retrieve the best outlier by an objective function on a calibration set with a very large sample size. However, the training process of reinforcement learning is not robust and is difficult to converge compared to other PTQ methods, making the calibration process very inefficient.
Large Language Model Quantization. Currently, large language model quantization is mainly divided into quantization-based training (Liu et al., 2023; Bondarenko et al., 2024; Zhu et al., 2023) and post-training quantization. Quantization during training is very difficult to apply due to the large amount of computational resources and overhead required, as well as the large amount of training data. The remaining methods such as ZeroQuant (Yao et al., 2022) utilize knowledge distillation to train one transformer layer at a time instead of training all transformer layers directly. However, poor quality of certain data can also affect the performance of the model after quantization (Wang et al., 2024a). Post-training quantization requires less data compared to quantization-aware training, and only a few thousand or even a few hundred calibration sets are needed to complete the retrieval of sensitive weights for the model. This greatly reduces the data cost and the computational cost of model quantization. Both GPTQ (Frantar et al., 2022a) and SPQR (Dettmers et al., 2023), quantization methods derived from the OBQ method, have shown superior performance in large language models. There are also methods such as AWQ (Lin et al., 2024), LLM.int8() (Dettmers et al., 2022) that determine the outlier of the model by searching for the activation of the model, and then quantify it with the corresponding method. However, the method of searching for outliers by activation has poor interpretability.
Table 1: Gradient quantization of non-outliers with Hessian in comparison to the C4 dataset (Raffel et al., 2020). We localize outliers by gradient and Hessian on the llama-2-hf series with the llama-3-hf model and quantify non-outliers with RTN (Nagel et al., 2020) to 4 bits.
FP16(%) base on Gradients FP16(%) base on Hessian
[…]
Llama-2-7b-hf 6.97 13.4328 9.3408 7.1161 13.4358 9.3442 7.1187
Llama-2-13b-hf 6.47 11.5711 7.7271 6.5120 11.5755 7.7299 6.5142
Llama-3-8b-hf 8.99 18.4334 14.4928 11.5442 18.4390 14.4942 11.5483
3 The Proposed Method
The outlier search method based on Hessian matrix (Frantar et al., 2022a; Dettmers et al., 2023) is built upon the OBQ approach (Frantar et al., 2022b). The OBQ method retrieves the parameters that have the least impact on the model parameters by using the following equation:
$\begin{split}\Delta E=\sum_{i}g_{i}\Delta w_{i}+\frac{1}{2}\sum_{i}h_{ii}\Delta w_{i}^{2}+\\ \frac{1}{2}\sum_{i\neq j}h_{ij}\Delta w_{i}\Delta w_{j}+O\left(\Delta w^{3}\right),\end{split}$ (1)
where $g_{i}=\frac{\partial E}{\partial w_{i}}$ is the first order gradient. In the OBQ framework, it is suggested that once a pre-trained model has fully converged, its gradients $g_{i}=\frac{\partial E}{\partial w_{i}}$ should ideally approach zero. However, our experiments demonstrate that LLMs continue to generate gradients in response to various text inputs. Bondarenko et al. (2023) denotes that the occurrence of outliers in large models arises from attention heads attempting to avoid updating hidden states. During this process, the softmax function magnifies the formation of strong outliers. Building on this observation, we hypothesize that when a well-trained LLM computes gradients for text input, it often focuses these gradients on irrelevant outliers as a mechanism to prevent hidden state updates. As shown in Fig. 1, compared to SPQR, GWQ demonstrates a sparser allocation of outliers. Specifically, the outliers identified are distributed either per-channel or per-row, which contrasts with the more dense and uniform distribution of outliers observed in SPQR. This sparsity in GWQ results in a more reasonable allocation, as it does not require to concentrate on the compensation of sensitive weight errors across every layer output. As shown in Tab. 1, with different proportions of outliers, the model using Hessian matrix search and quantized using RTN (Nagel et al., 2020) performs worse than the model using gradient search and quantized using RTN on the C4 dataset (Raffel et al., 2020), so it can be learned that it is more reasonable to use the gradient to do outliers localization.
Figure 1: Comparison of GWQ and SPQR outliers. Taking query (Q) in Self-Attention as an example, the outliers localized by GWQ (left figure) are more sparse compared to SPQR (right figure), and the row distribution is more pronounced, which is more friendly to subsequent quantization.
Figure 2: Sensitive weight Location. GWQ utilizes a single calibration sample to calibrate and backpropagates this sample with the subsequent token from the large model’s output as the target label. By halting the backpropagation process, GWQ is able to capture the model gradient $g$ prior to the update of the weights $W$. Within each layer, the gradients associated with each weight block are sorted in relation to their respective weights, and the top 1% with the largest $\left|g\right|$ are identified as the model’s outliers, which are the weights that are particularly sensitive.
3.1 Sensitive Weight Location
GWQ captures the absolute gradients with respect to LLM, acquired through back-propagation, the detail is shown in Alg. 1. We disabled bias in all selected linear layers, since the bias term does not get multiplied by inputs in the same way as weights do. This can cause the weight gradients to be unrepresentative of the actual importance of the weights, leading to a disproportionate influence on the MSE loss calculation. As it shown in Fig. 2, the calibration dataset is represented as $D_{c}$, the weights of LLM as $W$, the loss function of LLM as $\mathcal{L}=(W;D_{c})$. The gradients $g$ to the LLM is:
$g=\bigtriangledown_{W}\mathcal{L}=\bigtriangledown_{W}\mathcal{L}(W;D_{c}).$ (2)
In order to locate the sensitivities as outliers, we calculate loss $\mathcal{L}(W_{Q})$ after model quantization based on Eq. (3).
$\mathcal{L}(W_{Q})\simeq\mathcal{L}(W)-g^{\top}(W-W_{Q}).$ (3)
In order to search for quantitatively optimal outliers, we define the quantization process as an optimization problem whose process can be expressed as follows:
$Q\left(W\right)=arg\min_{Q}\left\|W-W_{Q}\right\|_{2}^{2}.$ (4)
Therefore the final weight optimization formula can be expressed as:
$Q\left(W\right)=arg\min_{Q}\bigtriangledown_{W}\mathcal{L}(W;D_{c})(W-W_{Q}).$ (5)
Algorithm 1 GWQ Gradient Accumulation
func get_gradient($M,X$)
1:a:=$a:=\emptyset$triangleright Set of accumulated gradients
[…]
3:  $o$ := M($X_{i}[0]$) $\triangleright$ model output
4:  $l$ := CrossEntropy($o$,$X_{i}[1]$)
5:  $l$.backward()
6:  $a$ := accum_grad($M$)
7:  $M$.zero_grad()
[…]
9:  $a_{j}:=\frac{a_{j}}{\text{dim}(X)}$
10:  $M_{g}:=a_{j}$triangleright update the model gradients
11:return $M$
func accum_grad($M,a$)
1:$q$ := [self_attn, mlp] //modules to be quantized
2:for $i=1,\dots,\text{dim}(a)$ do
3:  $a_{i}:=a_{i}+M_{a_{i}}$triangleright accumulate the gradients of each module in the model
4:return $a$
3.2 Gradient Aware Quantization
To quantize the LLM weight $W$, we first compute the scale $s$ and the zero point $z$, which are calculated as depicted in Eq. (6) and (7), where $\beta$ represents the group size (set to 16 in this case) and $b$ is the bit-width for quantization. The detail is shown in Alg. 2.
$s=\frac{\max(W)-\min(W)}{2^{b-1}},$ (6)
$z=\frac{-\min{W}}{s}.$ (7)
We select the weight by absolute values of the weight gradients $|W_{grad}|$, the largest 1% of them are filtered as outliers, and the rest are used as quantization weights. This process can be expressed as:
$O=\mathbb{I}\left(|W_{grad}|\geq\text{quantile}(\alpha_{1})\right),$ (8)
where $\alpha_{1}$ is the specific quantile value of the magnitude of weight gradients, which are used to identify outliers matrix $O$.
The quantization and dequantization operations are carried out per channel within each group to assess the quantization error. The formulas for quantization and dequantization are provided below (Eq. (9) and (10)), and any identified outliers are excluded from the error computation (Eq. (8)):
$q=round(\frac{W}{s})+z.$ (9)
The quantized weight $Q(W)$ can be denoted as:
$Q(W)=s\times(q-z).$ (10)
Algorithm 2 GWQ Quantization Algorithm
func GWQ_Quantize($M,X,b,\beta,p$)
Input: M – the model to be quantized
$X\in\mathbb{R}^{n\times d}$ – calibration data
$b$ – base number of quantization bits
$\beta$ – quantization group sizes
$p$ – global outlier percentage
1:$M$ :=get_gradient($M,X$)
2:$L$ :=$M$.layers
3:$n$ := dim(layers)
4:$l_{i_{s}}$ := $L_{i}$.sublayers
5:for $i=1,\dots,\text{dim}(L)$ do
6:for $j=1,\dots,\text{dim}(l_{i_{s}})$ do
7:  $W:=$
[…]
10:  $l_{i_{s}}.W:=W$
11:return M
func GWQ_Sublayer_Quantize($W,W_{g},b,\beta,p_{s}$)
Input: $W\in\mathbb{R}^{m\times n}$ – weight matrix
$W_{g}\in\mathbb{R}^{m\times n}$ – gradient of weight matrix
$p_{s}$ – sublayer outlier percentage
1:$Q$ := int_matrix(m,n) //temporary quantized weight
2:$O$ := $\mathbb{I}\left(|W_{g}|\geq\text{quantile}(\alpha_{1})\right)$
3:for $i=1,\beta 1,2\beta 1,\dots,n$ do
4:  $\vec{s},\vec{z}$ := fit_quantizer($W_{:,i:i+\beta}$, $O_{:,i:i+\beta}$)
5:  for $j=i,\dots,i+\beta$ do
6:   $Q_{:,j}$ := quantize($W_{:,j},\vec{s},\vec{z}$)
7:   $\vec{w_{q}}$ := dequantize($Q_{:,j},\vec{s},\vec{z}$)
8:   $W_{:,j}:=\vec{w_{q}}\dot{(}1-O_{:,j}))$
9:return $W$
func fit_quantizer($M,\beta$)
1:$m_{i}:=\text{flatten}(M)$
2:$s_{i}:=\text{vectors}(O)$
3:for $i=1,\beta_{1},\beta_{2},\dots,\text{dim}(m)$ do
4:  $s_{i}:=\frac{\text{max}(m)-\text{min}(m)}{\text{dim}(m)}$
5:  $z_{i}:=\min(\frac{m-s_{i}}{s_{i}})$
6:return $s,z$
func quantize($W,s,z$)
1:return $\lfloor\frac{W}{s}+z+0.5\rfloor$
func dequantize($Q,s,z$)
1:return $s\cdot(Q-z)$
4 Experiments
Table 2: Perplexity of Llama-2-hf family and Llama-3-8B-hf on language task datasets. GWQ outperforms other quantization methods, especially when the average bit-width of the model is significantly lower.
Llama-2-7B-hf Llama2-13B-hf Llama-3-8B-hf
[…]
Original 16 - 5.47 6.97 4.88 6.47 6.23 8.99
+ GPTQ 4 1024 5.83 7.79 5.12 7.72 8.21 11.48
+ AWQ 4 512 5.60 7.70 4.97 7.10 6.64 10.69
+ SPQR 4.63 1024 5.53 7.02 4.93 6.50 6.42 9.23
+ SPQR 3.98 1024 5.55 7.06 4.95 6.53 6.45 9.26
+ GWQ-O (Ours) 4.63 1 5.51 7.01 4.91 6.48 6.40 9.21
+ GWQ-R (Ours) 3.98 1 5.53 7.04 4.94 6.52 6.43 9.25
Table 3: Performance of the Llama-2-hf family and Llama-3-8B-hf on zero-shot tasks on language task datasets. GWQ demonstrates a lower loss of accuracy compared to other quantization methods.
Llama-2-7B-hf Llama-2-13B-hf Llama-3-8B-hf
Acc.(%$\uparrow$ Avg_bit cab_num WikiText2 C4 WikiText2 C4 WikiText2 C4
Original 16 - 61.01 56.77 62.88 57.90 58.34 51.71
+ GPTQ 4 1024 59.88 54.74 61.91 56.18 43.78 47.75
+ AWQ 4 512 60.54 55.16 62.53 56.44 57.07 48.96
+ SPQR 4.63 1024 60.80 56.63 62.71 57.78 57.70 51.20
+ SPQR 3.98 1024 60.56 55.18 62.55 56.45 57.09 48.97
+ GWQ-O (Ours) 4.63 1 60.83 56.65 62.73 57.80 57.71 51.22
+ GWQ-R (Ours) 3.98 1 60.58 55.20 62.56 56.47 57.11 48.99
Overview. First, we validate the superiority of our method over other quantization approaches on the Llama-2 family (Touvron et al., 2023b), providing detailed compression metrics for comparison. Additionally, we present the results of our method compressed to 2 bits versus 3 bits, confirming its reliability at lower bit precision. Following this, we demonstrate the quantization performance of our method on other models, showcasing its generalizability. Finally, we present zero-shot quantization results on multimodal models to verify the multimodal context generalizability of our approach.
4.1 Experiments Setting
Setups. The language-only models we selected include Llama2-7B and 13B (Touvron et al., 2023b), Falcon-7B (Penedo et al., 2023), and Mistral-7B-v0.1 (Jiang et al., 2023). These models were quantized using the first sample of the RedPajama dataset. To evaluate Perplexity (PPL) and Accuracy (Acc.), we utilize the WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020) validation sets. All experiments were conducted on an NVIDIA A100 (80G) GPU.
Baselines. Our primary baselines include AWQ (Lin et al., 2024) and GPTQ (Frantar et al., 2022a), while we benchmark our approach against the state-of-the-arts SPQR (Dettmers et al., 2023). To ensure that the average bit-width of the model remains below 4 bits—surpassing GPTQ and AWQ in compression—we quantize the non-outlier weights at 4-bit and 3-bit precision. These configurations are denoted GWQ-O for 4-bit quantization, GWQ-R for 3-bit quantization.
4.2 Main Result
We conduct separate evaluations for both pure large language models and large multimodal models. For the large language models, we assess the accuracy of the quantized model in terms of perplexity, zero-shot tasks, and quantized inference memory overhead and latency. In addition to evaluating performance on linguistic tasks for the large multimodal models, we also test the zero-shot target detection task across multiple datasets.
Table 4: Perplexity of Qwen-VL family models on language tasks datasets. GWQ quantized model has a lower PPL compared to other current quantization methods, and the PPL also remains essentially constant at lower bits.
QWEN-VL QWEN-VL-chat
[…]
Original 16 8.21 10.10 9.85 11.87
+ GPTQ 4 9.31 11.97 11.50 13.59
+ AWQ 4 8.69 10.98 11.02 12.88
+ SPQR 4.63 8.54 10.25 10.13 12.02
+ GWQ-O (Ours) 4.63 8.24 10.10 9.86 11.88
+ GWQ-R (Ours) 3.63 8.53 10.42 10.94 12.62
Table 5: Performance of the Qwen-VL family on the zero-shot tasks on language tasks datasets. GWQ has lower loss of accuracy compared to other quantization methods.
QWEN-VL QWEN-VL-chat
Acc.(%$\uparrow$ Avg_bit WikiText2 C4 WikiText2 C4
Original 16 54.18 49.51 52.06 48.07
+ GPTQ 4 52.07 47.08 49.83 46.09
+ AWQ 4 53.58 48.12 50.77 47.43
+ SPQR 4.63 54.05 49.26 51.99 47.35
+ GWQ-O (Ours) 4.63 54.06 49.50 52.03 48.07
+ GWQ-R (Ours) 3.63 53.87 49.32 52.01 49.98
Table 6: Performance of the Qwen-VL family on the zero-shot tasks on target detection datasets. In the three RefCOCO datasets, GWQ is more accurate compared to SPQR for the same bits, achieving less quantization loss.
QWEN-VL QWEN-VL-chat
Acc.(%$\uparrow$ Avg_bit test-A test-B val test-A test-B val
Original 16 92.26 85.34 89.36 92.27 84.51 88.55
+ GPTQ 4 90.61 83.14 87.13 90.33 82.27 86.67
+ AWQ 4 91.33 84.51 88.01 91.14 83.61 87.77
+ SPQR 4.63 92.17 84.87 88.97 91.76 84.20 88.48
+ SPQR 3.98 91.33 84.51 88.01 91.15 83.61 87.77
+ GWQ-O (Ours) 4.63 92.18 84.88 88.98 91.78 84.22 88.49
+ GWQ-R (Ours) 3.98 91.34 84.52 88.03 91.16 83.62 87.79
Table 7: Comparison of speedup and memory consumption of different quantization methods. GWQ significantly improves the inference speed and reduces the memory required compared to the original model.
Llama2-7B-hf Llama2-13B-hf
Methods Avg_bit Speedup($\uparrow$) Memory($\downarrow$) Speedup($\uparrow$) Memory($\downarrow$)
Original 16 $\times$1 12.8 $\times$1 23.6
+ GPTQ 4 $\times$2.0 3.1 $\times$2.0 6.7
+ AWQ 4 $\times$2.1 3.1 $\times$2.1 6.7
+ SPQR 4.63 $\times$1.2 N/A $\times$1.2 N/A
+ GWQ (Ours) 4.63 $\times$1.2 4.1 $\times$1.2 7.1
Language Model. As shown in Tab. 2 and Tab. 3, we first evaluate the perplexity of the quantized model on the WikiText2 Merity et al. (2016) and C4 Raffel et al. (2020) datasets on the Llama2-7B and 13B models. Additionally, we assess the models’ ability to zero-shot on these three datasets and compare the results to several baseline methods. GWQ achieves a lower average bit-width than other quantization methods by applying more aggressive compression to the non-outlier weights. Compared to other quantization approaches, GWQ yields a lower PPL on all datasets under identical bit compression. For the zero-shot task, the Acc. of the quantized model of GWQ is significantly higher than that of other quantization methods. indicating that GWQ results in less performance degradation. Notably, under extreme compression, GWQ sustains comparable performance to 4-bit quantization, even at 3-bit.
Figure 3: Performance of Different Quantized Qwen-VL. GWQ performs well compared to other quantization methods in the zero-shot grounding detection task.
Table 8: Performance of GWQ at 4bit quantization for different number of calibration samples. The quantization loss of the model essentially does not vary with the number of different calibration samples, allowing for low-loss quantization with fewer samples.
num_sample=1 num_sample=4 num_sample=8 num_sample=12
[…]
Llama2-7B-hf 5.51 7.01 5.50 7.01 5.51 7.00 5.50 6.99
Llama2-13B-hf 4.91 6.48 4.90 6.49 4.91 6.47 4.91 6.48
Llama-3-8B-hf 6.40 9.21 6.39 9.20 6.38 9.20 6.40 9.20
Table 9: Performance of GWQ at 4bit quantization for different sample of calibration samples. The model quantization loss is essentially invariant to changes in the quantization sample and is less dependent on sample type.
sample one sample two sample three sample four
[…]
Llama2-7B-hf 5.51 7.01 4.91 6.97 5.47 6.97 5.47 6.97
Llama2-13B-hf 4.91 6.48 4.88 6.47 4.88 6.47 4.88 6.47
Llama-3-8B-hf 6.40 9.21 5.25 7.75 5.25 7.75 5.25 7.75
Multimodal Models. For the multimodal model, we evaluate the performance of the compressed model in language tasks and multimodal tasks with the Qwen-VL family (Bai et al., 2023a, b). Similar to the language models, we evaluate the language generation capabilities of the quantized model on the WikiText2, C4 dataset. The multimodal tasks are mainly focused on zero-shot target detection. We choose the RefCOCO (Kazemzadeh et al., 2014; Yu et al., 2016) dataset to evaluate the zero-shot capability of the quantized model compared to other quantization methods. Intersection over Union $IoU$, which is denoted as:
$IoU=\frac{Box_{a}\cap Box_{b}}{Box_{a}\cup Box_{b}},$ (11)
where $Box_{b}$ denotes the human-labeled correct result in the dataset labeling (Ground-Truth Box) and $Box_{a}$ denotes the result predicted by the algorithm (Predicted Box).
In the experiment setting, the threshold of $IoU$ is 0.5, so the target detection accuracy Acc. can be defined as:
$Acc.=\frac{correct\_num}{total\_num},$ (12)
where $correct\_num$ is the number of samples detected correctly and $total\_num$ is the total number of samples.
As it shown in Tab. 4 and Tab. 5, quantized qwen-vl family model of GWQ outperforms the current stat-of-the-arts in both the language task and the zero-shot test, showing better performance. The GWQ quantized MLLM performs better on the verbal task compared to other quantization methods, so we will use the quantized model to test it on the visual grounding task. As it shown in Tab. 6, in the zero-shot grounding detection task, both GWQ-quantified Qwen-VL family models with currently outperform the current stat-of-the-arts on most of the datasets, achieving a generalization of the multimodal model, The performance is shown in Fig. 3, for the task of visual grounding detection of long context, the GWQ post-quantization model is more advantageous than SPQR post-quantization.
Model Efficiency Comparison. For the model throughput (tokens/s), the speedup rate can be obtained by calculating the time required to generate 512 tokens after quantization and comparing it with the original model. For the model inference overhead (G), we use the peak value when the model generates 512 tokens as the maximum inference memory by recording it.
As illustrated in Tab. 7, the GWQ quantized model can achieve about 1.2$\times$ speedup relative to the original model, which is on par with the current mixed-accuracy method SPQR. Compared to the full-precision quantization method, GWQ results in higher inference latency due to floating-point conversion during inference. Compared to the original model, GWQ consumes less memory during inference, which makes it easier to deploy on edge devices.
Figure 4: Effect of calibration sample on quantification. For different numbers of calibration sets, the PPL of the quantized model tends to be essentially constant on WikiText2 and C4. Therefore, the effect of the number of calibration set on the quantization effect is negligible.
4.3 Ablation Study
In this section, we examine the impact of varying the number of calibration samples on the performance of our method. Tab. 8 presents the results for the calibration sample sizes 1, 4, 8 and 12, evaluated on the WikiText2 and C4 datasets. Tab. 8 and Fig. 4 indicate that increasing the number of calibration samples has minimal impact on the PPL of the quantized model. This suggests that the size of the calibration set has little influence on the quantization performance of GWQ. Compared to most of the current quantization methods, which require large datasets, GWQ has a lower calibration cost and can achieve calibration with fewer samples.
To further investigate the impact of different calibration samples, we selected four distinct samples, each of which was used as a calibration set to quantize the model. According to Tab. 9, when different calibration samples are used to quantize the same model, the resulting PPL values on the same dataset remain similar, indicating that the choice of calibration samples has a negligible effect on the quantization results. Unlike AWQ, GWQ can be well adapted to the calibration sets of different samples, and the quantized model performs well even if the calibration sets differ greatly.
5 Conclusion
This paper introduces GWQ, a weight-only post-training mixed-precision quantization approach based on first-order gradients. GWQ preserves the weights corresponding to the top 1% of outliers at FP16 precision, while storing the remaining non-outlier weights in a lower-bit format. GWQ found that it is easier to obtain more accurate outliers using the first-order gradient to locate the sensitive weights of the pre-trained model than the current search for model outliers via the Hessian matrix. GWQ achieves the lowest quantization loss compared to current state-of-the-arts methods by utilizing only a single calibration sample. Considerable results are also obtained by GWQ on multimodal models, which can achieve better results on most of the zero-shot target detection tasks compared to the current stat-of-the-art methods. GWQ achieves 1.2$\times$ inference acceleration compared to the original model internship and uses less memory in the inference process.
6 Limitation
GWQ gets gradients via back propagation and ranks them accordingly. However, in models such as OPT, the activation function is ReLU, and the correct gradients cannot be computed because of gradient vanishing. Additionally, backpropagation demands substantial memory resources, a single A100 80G GPU is insufficient to handle the large computational requirements. Furthermore, since GWQ employs mixed-precision quantization, it is less hardware-friendly compared to methods like AWQ, resulting in higher inference latency after quantization. We will continue to refine and optimize the GWQ methodology.

### v3

@@
-GWQ: Gradient-Aware Weight Quantization for Large Language Models
+Gradient-Aware Weight Quantization for Large Language Models
 Abstract
 Large language models (LLMs) show impressive performance in solving complex language tasks.
-However, its large number of parameters present significant challenges for the deployment and application of the model on edge devices.
-Compressing large language models to low bits can enable them to run on resource-constrained devices, often leading to performance degradation.
+However, its large number of parameters presents significant challenges for the deployment.
+So, compressing LLMs to low bits can enable to deploy on resource-constrained devices.
 To address this problem, we propose gradient-aware weight quantization (GWQ), the first quantization approach for low-bit weight quantization that leverages gradients to localize outliers, requiring only a minimal amount of calibration data for outlier detection.
-GWQ retains the weights corresponding to the top 1% outliers preferentially at FP16 precision, while the remaining non-outlier weights are stored in a low-bit format.
-GWQ found experimentally that utilizing the sensitive weights in the gradient localization model is more scientific compared to utilizing the sensitive weights in the Hessian matrix localization model.
-Compared to current quantization methods, GWQ can be applied to multiple language models and achieves lower PPL on the WikiText2 and C4 dataset.
-In the zero-shot task, GWQ quantized models have higher accuracy compared to other quantization methods.
-GWQ is also suitable for multimodal model quantization, and the quantized Qwen-VL family model is more accurate than other methods.
-Zero-shot target detection task dataset RefCOCO outperforms the current stat-of-the-arts method SPQR.
-GWQ achieves 1.2$\times$ inference speedup in comparison to the original model, and effectively reduces the inference memory.
-GWQ: Gradient-Aware Weight Quantization for Large Language Models
-Yihua Shao1,2, Siyu Liang3, Zijian Ling3, Minxi Yan3, Haiyang Liu4, Siyu Chen2, Ziyang Yan5, Chenyu Zhang5, Haotong Qin6*, Michele Magno6, Yang Yang2, Zhen Lei2, Yan Wang3, Jingcai Guo7, Ling Shao8, Hao Tang1* 1PKU  2CASIA  3THU  4USTB  5UNITN  6ETHz  7PolyU  8UCAS thanks: The first two authors contributed equally.
+GWQ retains the top 1% outliers preferentially at FP16 precision, while the remaining non-outlier weights are stored in a low-bit.
+We widely evaluate GWQ on different task include language modeling, grounding detection, massive multitask language understanding and vision-language question and answering.
+Results show that models quantified by GWQ performs better than other quantization method.
+During quantization process, GWQ only need one calibration set to realize effective quant.
+Also, GWQ achieves 1.2$\times$ inference speedup in comparison to the original model and effectively reduces the inference memory.
+Gradient-Aware Weight Quantization for Large Language Models
+Yihua Shao1,2, Yan Gu3, Siyu Chen3, Haiyang Liu3, Zijian Ling4, Minxi Yan2, Ziyang Yan5, Chenyu Zhang5, Michele Magno6, Haotong Qin6*, Yan Wang3, Jingcai Guo7, Ling Shao1* Hao Tang1* 1PKU  2CASIA  3THU  4USTB  5UNITN  6ETHz  7PolyU  8UCAS thanks: The first two authors contributed equally.
 correspondence authors: qinhaotong@gmail.com, haotang@pku.edu.cn
 1 Introduction
-Large language models (LLMs) (Touvron et al., 2023a; Achiam et al., 2023; Almazrouei et al., 2023; Touvron et al., 2023b) based on Transformer (Vaswani, 2017) have demonstrated their outstanding capabilities in scenarios such as complex question and answer, language modeling (Brown, 2020), and so on.
-It can often generate more accurate answers with different inputs.
-Its ability to handle complex linguistic tasks is due to its huge pre-trained corpus (Kaplan et al., 2020) and an astronomical number of parameters (Chowdhery et al., 2023).
-Its outstanding performance is rapidly generating huge applications in people’s lives.
-However, the huge memory consumption generated by its pre-training and inference phases leads to great difficulties in its application and deployment (Li et al., 2024a).
+Large language models (LLMs) (Touvron et al., 2023a; Achiam et al., 2023; Almazrouei et al., 2023; Touvron et al., 2023b) based on Transformer (Vaswani, 2017) have demonstrated their outstanding capabilities in scenarios such as language modeling (Brown, 2020).
+Its ability to handle complex linguistic tasks is due to its large pre-trained data (Kaplan et al., 2020; Yan et al., 2024, 2025) and an astronomical number of parameters (Chowdhery et al., 2023).
+However, the huge memory consumption leads to great difficulties in its application and deployment (Li et al., 2024a; Remondino et al., 2023).
 Therefore, when deploying large language models, there are often huge GPU clusters to support the inference of the models (Xu et al., 2024).
-In order to make large language models adaptable to resource-constrained edge devices (Li et al., 2024b; Huang et al., 2024), model compression (Ma et al., 2023; Gu et al., 2023) has become a common means to reduce the computational memory of models.
-Among them, post-training large language model quantization (Dettmers and Zettlemoyer, 2023) is a commonly used method to reduce the model inference overhead.
-Most current quantization methods compress the model to 3 or 4 bits (Dettmers et al., 2023; Frantar et al., 2022a), while trying to ensure that the model has as little loss of accuracy and performance as possible.
-Post-training quantization of pre-trained models reduces model inference overhead by allowing the number of parameter bits to decrease.
-This approach produces a loss of accuracy that affects the performance of the model in use.
-Therefore, to reduce the size of the model while ensuring its performance, it is necessary to design effective quantization algorithms (Lin et al., 2024).
-With effective quantization algorithms, researchers can make the performance of low-bit models infinitely close to that of 16-bit models.
-OBQ (Frantar and Alistarh, 2022) indicates that when the pre-trained model has fully converged, the model should exhibit zero gradients.
-However, our experiments reveal that the LLMs still generate gradients in response to different text inputs, indicating that the model has not completely converged and remains sensitive to the input.
+In order to apply LLMs to resource-constrained edge devices (Li et al., 2024b; Huang et al., 2024), model compression (Ma et al., 2023; Gu et al., 2023) has become common means to reduce the computational memory.
+Among them, post-training quantinazation of LLMs (Dettmers and Zettlemoyer, 2023) is a widely used method.
+Most current quantization methods target to compress modela to 3 or 4 bits (Dettmers et al., 2023; Frantar et al., 2022a), while ensuring models has less loss of performance.
+Therefore, reducing the size of the model while ensuring its performance is necessary by design effective quantization algorithms (Lin et al., 2024b).
+With effective quantization algorithms, researchers can make low-bit models’ performance close to that of 16-bit models.
+OBQ (Frantar and Alistarh, 2022) indicates that when the pre-trained model has converged, the model should exhibit zero gradients.
+However, our experiments reveal that the LLMs still generate gradients in response to different text inputs.
 Inspired by this phenomenon, we propose a gradient-aware post-training weight-only quantization method called GWQ.
 GWQ is the first post-training quantization approach to utilize gradients to locate outliers in pre-trained models.
-It introduces two key innovations.
-First, GWQ selects 1% of the weights as outliers by looking at the magnitude of the gradient for the response to the calibration set, preserving these outliers in FP16 precision while quantizing the remaining 99% of the weights to 4 bits or 3 bits.
+As same as other quantization method Lin et al.
+(2024a); Dettmers et al.
+(2023), GWQ selects 1% sensitive weights as outliers by searching the gradient for the response to the calibration set, preserving these outliers in FP16 precision while quanting the remaining 99% of the weights to 4 bits or 3 bits.
 In summary, the primary contributions of this paper are as follows:
   • 1)
-    GWQ discovered that employing the first-order gradient to search for and locate sensitive weights is more rational than using the Hessian matrix for localization, and the model loss after quantization is also lower.
-Therefore, using a first-order gradient to locate outliers is more reasonable.
+    GWQ discovered that locating sensitive weights by first-order gradient is more rational than Hessian matrix.
   • 2)
-    GWQ is the first accurate first-order gradient-aware post-training weight quantization method for pre-trained LLMs, requiring only a minimal quantity of calibration data to identify outliers efficiently.
+    GWQ is the first accurate first-order gradient-aware post-training weight quantization method for pre-trained LLMs, requiring only one of calibration data to locating outliers.
   • 3)
-    GWQ outshines the current state-of-the-arts method SPQR on the wikitext and C4 datasets.
-The majority of zero-shot detection tasks in the multimodal dataset RefCOCO have reached the state-of-the-arts.
-Meanwhile, the quantified model has achieved a 1.2$\times$ acceleration compared to the original model and utilized less memory during inference.
+    GWQ achieves state-of-the-art (SOTA) in language modeling, grounding detection, massive multitask language understanding and vision-language question and answering tasks in different LLMs.
+Meanwhile, the quantified models have achieved 1.2$\times$ acceleration compared to the original model and cost less memory during inference.
 2 Related Work
@@
 Post-training quantization (PTQ) applies mainly to visual models (Gholami et al., 2022).
-For example, AdaRound (Nagel et al., 2020), AdaQuant (Hubara et al., 2020), BitSplit (Wang et al., 2024b) can be used for models with a small number of parameters.
-When models have a large number of parameters, they cause a significant loss of accuracy, and the reduction in memory overhead of the quantized model compared to the original model is not particularly significant.
-Most current quantization methods for models are improved by Optimal Brain Quantization (OBQ) (Frantar and Alistarh, 2022).
-OBQ is improved by the Optimal Brain Surgeon (OBS) (Hassibi et al., 1993; Singh and Alistarh, 2020; Frantar et al., 2021), both OBS and OBQ default to a model whose response to input is 0 at the end of training, but the OBQ method applies the idea of weight pruning to model quantization by quantifying the weights that have the least impact on the network and compensating for them with the remaining weights.
-Some current PTQ methods perform outlier search by reinforcement learning (Wang et al., 2024b) to retrieve the best outlier by an objective function on a calibration set with a very large sample size.
-However, the training process of reinforcement learning is not robust and is difficult to converge compared to other PTQ methods, making the calibration process very inefficient.
+For example, AdaRound (Nagel et al., 2020), AdaQuant (Hubara et al., 2020), and BitSplit (Wang et al., 2024b) can be used for models with less parameters.
+When models have a large number of parameters, they cause a significant loss but the reduction in memory is not significant.
+Most quantization methods are developed by Optimal Brain Quantization (OBQ) (Frantar and Alistarh, 2022).
+OBQ is developed by the Optimal Brain Surgeon (OBS) (Hassibi et al., 1993; Singh and Alistarh, 2020; Frantar et al., 2021), which default a model response to input is 0 at the end of training.
+But the OBQ applies weight pruning to model quantization by quantifying the weights that have the less impact on the network.
+Some PTQ methods locating outlier by reinforcement learning (Wang et al., 2024b) to retrieve the best outlier on a calibration set with a very large sample size.
+However, the reinforcement learning is not robust and is difficult to converge compared to other PTQ methods, making the calibration process inefficient.
 Large Language Model Quantization.
-Currently, large language model quantization is mainly divided into quantization-based training (Liu et al., 2023; Bondarenko et al., 2024; Zhu et al., 2023) and post-training quantization.
+Currently, large language model quantization is mainly divided into quantization-based training (Liu et al., 2023; Bondarenko et al., 2024; Zhu et al., 2023; Shao et al., 2025b, a) and post-training quantization.
 Quantization during training is very difficult to apply due to the large amount of computational resources and overhead required, as well as the large amount of training data.
@@
 Both GPTQ (Frantar et al., 2022a) and SPQR (Dettmers et al., 2023), quantization methods derived from the OBQ method, have shown superior performance in large language models.
-There are also methods such as AWQ (Lin et al., 2024), LLM.int8() (Dettmers et al., 2022) that determine the outlier of the model by searching for the activation of the model, and then quantify it with the corresponding method.
+There are also methods such as AWQ (Lin et al., 2024b), LLM.int8() (Dettmers et al., 2022) that determine the outlier of the model by searching for the activation of the model, and then quantify it with the corresponding method.
 However, the method of searching for outliers by activation has poor interpretability.
-Table 1: Gradient quantization of non-outliers with Hessian in comparison to the C4 dataset (Raffel et al., 2020).
-We localize outliers by gradient and Hessian on the llama-2-hf series with the llama-3-hf model and quantify non-outliers with RTN (Nagel et al., 2020) to 4 bits.
+Figure 1: Compared of GWQ and SPQR. Compared to SPQR, the outliers searched by GWQ are much sparser and easier to quantization.
+Table 1: Gradient quantization of non-outliers with Hessian in comparison on C4 dataset (Raffel et al., 2020).
+We localize outliers by gradient and Hessian and quant the non-outliers with RTN (Nagel et al., 2020) to 4 bits.
 FP16 (%) base on Gradients FP16 (%) base on Hessian
 […]
-Llama-2-7b-hf 6.97 13.4328 9.3408 7.1161 13.4358 9.3442 7.1187
-Llama-2-13b-hf 6.47 11.5711 7.7271 6.5120 11.5755 7.7299 6.5142
-Llama-3-8b-hf 8.99 18.4334 14.4928 11.5442 18.4390 14.4942 11.5483
+Llama-2-7b-hf 6.9759 20.5896 16.8329 15.4319 24.0418 20.2053 18.5902
+Llama-2-13b-hf 6.4734 20.3549 16.6689 15.1590 23.9322 20.1040 18.3405
+Llama-3-8b-hf 8.9990 22.4648 18.9843 17.5901 22.4423 22.9650 19.3283
 3 The Proposed Method
 The outlier search method based on Hessian matrix (Frantar et al., 2022a; Dettmers et al., 2023) is built upon the OBQ approach (Frantar et al., 2022b).
-The OBQ method retrieves the parameters that have the least impact on the model parameters by using the following equation:
+The OBQ method retrieves the parameters that have the least impact on the model parameters could be expressed as Eq.
+(1):
 $\begin{split}\Delta E=\sum_{i}g_{i}\Delta w_{i}+\frac{1}{2}\sum_{i}h_{ii}\Delta w_{i}^{2}+\\ \frac{1}{2}\sum_{i\neq j}h_{ij}\Delta w_{i}\Delta w_{j}+O\left(\Delta w^{3}\right),\end{split}$ (1)
@@
 1, compared to SPQR, GWQ demonstrates a sparser allocation of outliers.
-Specifically, the outliers identified are distributed either per-channel or per-row, which contrasts with the more dense and uniform distribution of outliers observed in SPQR.
-This sparsity in GWQ results in a more reasonable allocation, as it does not require to concentrate on the compensation of sensitive weight errors across every layer output.
+Specifically, the outliers identified are distributed either per channel or per row, which contrasts with the more dense and uniform distribution of outliers observed in SPQR.
+This sparsity in GWQ results in a more reasonable allocation, as it does not require concentrating on the compensation of sensitive weight errors across every layer output.
 As shown in Tab.
 1, with different proportions of outliers, the model using Hessian matrix search and quantized using RTN (Nagel et al., 2020) performs worse than the model using gradient search and quantized using RTN on the C4 dataset (Raffel et al., 2020), so it can be learned that it is more reasonable to use the gradient to do outliers localization.
-Figure 1: Comparison of GWQ and SPQR outliers.
-Taking query (Q) in Self-Attention as an example, the outliers localized by GWQ (left figure) are more sparse compared to SPQR (right figure), and the row distribution is more pronounced, which is more friendly to subsequent quantization.
 Figure 2: Sensitive weight Location.
@@
 GWQ captures the absolute gradients with respect to LLM, acquired through back-propagation, the detail is shown in Alg.
-1.
-We disabled bias in all selected linear layers, since the bias term does not get multiplied by inputs in the same way as weights do.
+LABEL:gradient_alg.
+We disabled bias in all selected linear layers since the bias term does not get multiplied by inputs in the same way as weights do.
 This can cause the weight gradients to be unrepresentative of the actual importance of the weights, leading to a disproportionate influence on the MSE loss calculation.
 As it shown in Fig.
-2, the calibration dataset is represented as $D_{c}$, the weights of LLM as $W$, the loss function of LLM as $\mathcal{L}=(W;D_{c})$.
-The gradients $g$ to the LLM is:
-$g=\bigtriangledown_{W}\mathcal{L}=\bigtriangledown_{W}\mathcal{L}(W;D_{c}).$ (2)
+2, the calibration dataset is represented as $D_{c}$, the weights of LLM as $W$, and the loss function of LLM as $\mathcal{L}=(W;D_{c})$.
+So, the gradients $g$ of the LLM could be expressed as (2),
+$g=\bigtriangledown_{W}\mathcal{L}=\bigtriangledown_{W}\mathcal{L}(W;D_{c}),$ (2)
 In order to locate the sensitivities as outliers, we calculate loss $\mathcal{L}(W_{Q})$ after model quantization based on Eq. (3).
 $\mathcal{L}(W_{Q})\simeq\mathcal{L}(W)-g^{\top}(W-W_{Q}).$ (3)
-In order to search for quantitatively optimal outliers, we define the quantization process as an optimization problem whose process can be expressed as follows:
-$Q\left(W\right)=arg\min_{Q}\left\|W-W_{Q}\right\|_{2}^{2}.$ (4)
-Therefore the final weight optimization formula can be expressed as:
+We define the quantization process as an optimization problem to search for quantitatively optimal outliers. The process can be expressed as Eq. (4),
+$Q\left(W\right)=arg\min_{Q}\left\|W-W_{Q}\right\|_{2}^{2},$ (4)
+Therefore the final weight optimization formula can be expressed as Eq (5),
 $Q\left(W\right)=arg\min_{Q}\bigtriangledown_{W}\mathcal{L}(W;D_{c})(W-W_{Q}).$ (5)
-Algorithm 1 GWQ Gradient Accumulation
-func get_gradient($M,X$)
-1:a:=$a:=\emptyset$triangleright Set of accumulated gradients
-[…]
-3:  $o$ := M($X_{i}[0]$) $\triangleright$ model output
-4:  $l$ := CrossEntropy($o$,$X_{i}[1]$)
-5:  $l$.backward()
-6:  $a$ := accum_grad($M$)
-7:  $M$.zero_grad()
-[…]
-9:  $a_{j}:=\frac{a_{j}}{\text{dim}(X)}$
-10:  $M_{g}:=a_{j}$triangleright update the model gradients
-11:return $M$
-func accum_grad($M,a$)
-1:$q$ := [self_attn, mlp] //modules to be quantized
-2:for $i=1,\dots,\text{dim}(a)$ do
-3:  $a_{i}:=a_{i}+M_{a_{i}}$triangleright accumulate the gradients of each module in the model
-4:return $a$
 3.2 Gradient Aware Quantization
@@
 (6) and (7), where $\beta$ represents the group size (set to 16 in this case) and $b$ is the bit-width for quantization.
-The detail is shown in Alg.
-2.
 $s=\frac{\max(W)-\min(W)}{2^{b-1}},$ (6)
@@
 We select the weight by absolute values of the weight gradients $|W_{grad}|$, the largest 1% of them are filtered as outliers, and the rest are used as quantization weights.
-This process can be expressed as:
+The process can be expressed as Eq.
+(8),
 $O=\mathbb{I}\left(|W_{grad}|\geq\text{quantile}(\alpha_{1})\right),$ (8)
@@
 The formulas for quantization and dequantization are provided below (Eq.
-(9) and (10)), and any identified outliers are excluded from the error computation (Eq.
-(8)):
-$q=round(\frac{W}{s})+z.$ (9)
-The quantized weight $Q(W)$ can be denoted as:
+(9) and (10)), and any identified outliers are excluded from the error computation Eq.
+(8):
+$q=round(\frac{W}{s})+z,$ (9)
+Finally, the quantized weight $Q(W)$ can be denoted as Eq (10):
 $Q(W)=s\times(q-z).$ (10)
-Algorithm 2 GWQ Quantization Algorithm
-func GWQ_Quantize($M,X,b,\beta,p$)
-Input: M – the model to be quantized
-$X\in\mathbb{R}^{n\times d}$ – calibration data
-$b$ – base number of quantization bits
-$\beta$ – quantization group sizes
-$p$ – global outlier percentage
-1:$M$ :=get_gradient($M,X$)
-2:$L$ :=$M$.layers
-3:$n$ := dim(layers)
-4:$l_{i_{s}}$ := $L_{i}$.sublayers
-5:for $i=1,\dots,\text{dim}(L)$ do
-6:for $j=1,\dots,\text{dim}(l_{i_{s}})$ do
-7:  $W:=$
+4 Experiments
+Table 2: Results of Llama family on language modeling task. GWQ outperforms other quantization methods, especially when the average bit-width of the model is significantly lower.
+Average Calibration Llama-2-7B-hf Llama-2-13B-hf Llama-3-8B-hf
 […]
+Original 16 - 5.47 / 61.01 6.97 / 56.77 4.88 / 62.88 6.47 / 57.90 6.23 / 58.34 8.99 / 51.71
++ OPTQ 4 1024 9.76 / 51.53 11.24 / 49.01 9.37/ 51.89 10.65 / 50.02 11.56 / 39.34 15.84 /44.93
++ QuIP 4 1024 8.43 / 53.13 10.01 / 50.82 8.13 / 53.48 9.49 / 51.33 10.89 / 40.84 14.03 / 46.22
++ GPTQ 4 1024 5.83 / 59.88 7.79 / 54.74 5.12 / 61.91 7.72 / 56.18 8.21 / 43.78 11.48 / 47.75
++ AWQ 4 512 5.60 / 60.54 7.70 / 55.16 4.97 / 62.53 7.10 / 56.44 6.64 / 57.07 10.69 / 47.75
++ SPQR-O 4.63 1024 5.53 / 60.80 7.02 / 56.63 4.93 / 62.71 6.50 / 57.78 6.42 / 57.70 9.23 / 51.20
++ SPQR-R 3.98 1024 5.55 / 60.56 7.06 / 55.18 4.95 / 62.55 6.53 / 56.45 6.45 / 57.09 9.26 / 48.97
++ GWQ-O (Ours) 4.63 1 5.48 / 60.87 6.99 / 56.72 4.90 / 62.76 6.48 / 53.84 6.40 / 57.83 9.11 / 51.46
++ GWQ-R (Ours) 3.98 1 5.50 / 60.59 7.01 / 56.20 4.38 / 62.56 6.50 / 56.77 6.41 / 57.69 9.15 / 49.89
+Table 3: Results of QWEN-VL families on language molding task and grounding detection task. GWQ outperforms other quantization methods on both language modeling task and grounding detection task.
+Average Calibration QWEN-VL QWEN-VL-chat
 […]
-10:  $l_{i_{s}}.W:=W$
-11:return M
-func GWQ_Sublayer_Quantize($W,W_{g},b,\beta,p_{s}$)
-Input: $W\in\mathbb{R}^{m\times n}$ – weight matrix
-$W_{g}\in\mathbb{R}^{m\times n}$ – gradient of weight matrix
-$p_{s}$ – sublayer outlier percentage
-1:$Q$ := int_matrix(m,n) //temporary quantized weight
-2:$O$ := $\mathbb{I}\left(|W_{g}|\geq\text{quantile}(\alpha_{1})\right)$
-3:for $i=1,\beta 1,2\beta 1,\dots,n$ do
-4:  $\vec{s},\vec{z}$ := fit_quantizer($W_{:,i:i+\beta}$, $O_{:,i:i+\beta}$)
-5:  for $j=i,\dots,i+\beta$ do
-6:   $Q_{:,j}$ := quantize($W_{:,j},\vec{s},\vec{z}$)
-7:   $\vec{w_{q}}$ := dequantize($Q_{:,j},\vec{s},\vec{z}$)
-8:   $W_{:,j}:=\vec{w_{q}}\dot{(}1-O_{:,j}))$
-9:return $W$
-func fit_quantizer($M,\beta$)
-1:$m_{i}:=\text{flatten}(M)$
-2:$s_{i}:=\text{vectors}(O)$
-3:for $i=1,\beta_{1},\beta_{2},\dots,\text{dim}(m)$ do
-4:  $s_{i}:=\frac{\text{max}(m)-\text{min}(m)}{\text{dim}(m)}$
-5:  $z_{i}:=\min(\frac{m-s_{i}}{s_{i}})$
-6:return $s,z$
-func quantize($W,s,z$)
-1:return $\lfloor\frac{W}{s}+z+0.5\rfloor$
-func dequantize($Q,s,z$)
-1:return $s\cdot(Q-z)$
-4 Experiments
-Table 2: Perplexity of Llama-2-hf family and Llama-3-8B-hf on language task datasets.
-GWQ outperforms other quantization methods, especially when the average bit-width of the model is significantly lower.
-Llama-2-7B-hf Llama2-13B-hf Llama-3-8B-hf
-[…]
-Original 16 - 5.47 6.97 4.88 6.47 6.23 8.99
-+ GPTQ 4 1024 5.83 7.79 5.12 7.72 8.21 11.48
-+ AWQ 4 512 5.60 7.70 4.97 7.10 6.64 10.69
-+ SPQR 4.63 1024 5.53 7.02 4.93 6.50 6.42 9.23
-+ SPQR 3.98 1024 5.55 7.06 4.95 6.53 6.45 9.26
-+ GWQ-O (Ours) 4.63 1 5.51 7.01 4.91 6.48 6.40 9.21
-+ GWQ-R (Ours) 3.98 1 5.53 7.04 4.94 6.52 6.43 9.25
-Table 3: Performance of the Llama-2-hf family and Llama-3-8B-hf on zero-shot tasks on language task datasets. GWQ demonstrates a lower loss of accuracy compared to other quantization methods.
-Llama-2-7B-hf Llama-2-13B-hf Llama-3-8B-hf
-Acc.(%$\uparrow$ Avg_bit cab_num WikiText2 C4 WikiText2 C4 WikiText2 C4
-Original 16 - 61.01 56.77 62.88 57.90 58.34 51.71
-+ GPTQ 4 1024 59.88 54.74 61.91 56.18 43.78 47.75
-+ AWQ 4 512 60.54 55.16 62.53 56.44 57.07 48.96
-+ SPQR 4.63 1024 60.80 56.63 62.71 57.78 57.70 51.20
-+ SPQR 3.98 1024 60.56 55.18 62.55 56.45 57.09 48.97
-+ GWQ-O (Ours) 4.63 1 60.83 56.65 62.73 57.80 57.71 51.22
-+ GWQ-R (Ours) 3.98 1 60.58 55.20 62.56 56.47 57.11 48.99
+Original 16 - 8.21 / 54.18 10.10 / 49.51 89.36 9.85 / 52.06 11.87 / 48.07 88.55
++ OPTQ 4 1024 9.34 / 53.13 12.99 /45.76 85.78 13.78 / 50.02 15.71 / 44.93 84.12
++ QuIP 4 1024 9.48 / 53.87 12.68 / 46.41 86.92 12.33 / 51.65 13.98 / 45.76 85.41
++ GPTQ 4 1024 9.31 / 52.07 11.97 / 47.08 87.13 11.50 / 52.06 13.59 / 46.09 86.67
++ AWQ 4 512 8.69 / 53.58 10.98 / 48.12 84.51 11.02 / 50.77 12.88 / 47.43 87.77
++ SPQR-O 4.63 1024 8.54 / 54.05 10.25 / 49.26 88.47 10.13 / 51.99 12.02 / 47.35 88.26
++ SPQR-R 3.98 1024 8.69 / 53.57 10.98 / 48.14 88.00 11.03 / 50.80 12.88 / 46.84 87.78
++ GWQ-O 4.63 1 8.24 / 54.06 10.10 / 49.50 88.98 9.86 / 52.03 11.88 / 48.07 88.49
++ GWQ-R 3.98 1 8.53 / 53.87 10.01 / 49.12 88.42 10.34 / 51.93 12.62 / 47.18 87.93
+Figure 3: Performance of Different Quantized Qwen-VL. GWQ performs well compared to other quantization methods in the zero-shot grounding detection task.
 Overview.
-First, we validate the superiority of our method over other quantization approaches on the Llama-2 family (Touvron et al., 2023b), providing detailed compression metrics for comparison.
-Additionally, we present the results of our method compressed to 2 bits versus 3 bits, confirming its reliability at lower bit precision.
-Following this, we demonstrate the quantization performance of our method on other models, showcasing its generalizability.
-Finally, we present zero-shot quantization results on multimodal models to verify the multimodal context generalizability of our approach.
+We first apply GWQ to the Llama2 Touvron et al.
+(2023b) and Llama3 Dubey et al.
+(2024) family of models and evaluate its feasibility on a language modeling task.
+We then extend our approach to multimodal models Bai et al.
+(2023a, b); Liu et al.
+(2024), measuring the generalizability of our approach to multimodal scenarios with different base models for the grounding detection task, vision language question and answer (VQA).
+To ensure the fairness of the experiments, we compressed the average number of bits of all the mixed-precision quantized models to less than 4 bits.
 4.1 Experiments Setting
 Setups.
-The language-only models we selected include Llama2-7B and 13B (Touvron et al., 2023b), Falcon-7B (Penedo et al., 2023), and Mistral-7B-v0.1 (Jiang et al., 2023).
-These models were quantized using the first sample of the RedPajama dataset.
-To evaluate Perplexity (PPL) and Accuracy (Acc.), we utilize the WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020) validation sets.
-All experiments were conducted on an NVIDIA A100 (80G) GPU.
+The language-only models we selected include Llama2-7B and 13B (Touvron et al., 2023b) and Llama3-8b Dubey et al.
+(2024).
+As for vision-language model, we chose LLaVA-V1.5 Liu et al.
+(2024), Qwen-VL Bai et al.
+(2023a, b), Qwen-VL-chat Bai et al.
+(2023a, b).
+These models were quantized using the first sample of the RedPajama dataset Weber et al.
+(2024).
+To evaluate Perplexity (PPL) and Accuracy (Acc.), we utilize the WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020) as validation sets.
+All experiments were conducted on an NVIDIA A800 (80G) GPU.
+[…]
+$\text{PPL}=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P(x_{i}\mid x_{0:i-1})\right),$ (11)
+where $x_{i}$ represents the predicted token conditioned on the former tokens ($x_{<i}$) in inference step $i$ and $c_{w}$ denotes the average characters per word in the generated sequence.
+The next token prediction accuracy can be expressed as Eq. (18),
+$\mathrm{Acc}=\frac{N_{\hat{y_{t}}=\overrightarrow{y_{t}}}}{T},$ (12)
+where $\hat{y_{t}}$ is the predicted token by LLMs, $\overrightarrow{y_{t}}$ is the ground truth token, $T$ is the number of all tokens.
+The exact calculation of next token prediction accuracy will be explained in A.1.
 Baselines.
-Our primary baselines include AWQ (Lin et al., 2024) and GPTQ (Frantar et al., 2022a), while we benchmark our approach against the state-of-the-arts SPQR (Dettmers et al., 2023).
-To ensure that the average bit-width of the model remains below 4 bits—surpassing GPTQ and AWQ in compression—we quantize the non-outlier weights at 4-bit and 3-bit precision.
-These configurations are denoted GWQ-O for 4-bit quantization, GWQ-R for 3-bit quantization.
+Our baseline is categorized into uniform precision quantization and mixed precision quantization.
+For uniform precision quantization, we choose OPTQ Frantar et al.
+(2022b), QUIP Chee et al.
+(2023), GPTQ Frantar et al.
+(2022a) and AWQ Lin et al.
+(2024b) as the baseline.
+For mixed precision quantization, we choose SPQR Dettmers et al.
+(2023) as the baseline.
+For fairness, we test the results of SPQR and GWQ with an average bit less than 4 respectively.
 4.2 Main Result
@@
 In addition to evaluating performance on linguistic tasks for the large multimodal models, we also test the zero-shot target detection task across multiple datasets.
-Table 4: Perplexity of Qwen-VL family models on language tasks datasets.
-GWQ quantized model has a lower PPL compared to other current quantization methods, and the PPL also remains essentially constant at lower bits.
-QWEN-VL QWEN-VL-chat
-[…]
-Original 16 8.21 10.10 9.85 11.87
-+ GPTQ 4 9.31 11.97 11.50 13.59
-+ AWQ 4 8.69 10.98 11.02 12.88
-+ SPQR 4.63 8.54 10.25 10.13 12.02
-+ GWQ-O (Ours) 4.63 8.24 10.10 9.86 11.88
-+ GWQ-R (Ours) 3.63 8.53 10.42 10.94 12.62
-Table 5: Performance of the Qwen-VL family on the zero-shot tasks on language tasks datasets. GWQ has lower loss of accuracy compared to other quantization methods.
-QWEN-VL QWEN-VL-chat
-Acc.(%$\uparrow$ Avg_bit WikiText2 C4 WikiText2 C4
-Original 16 54.18 49.51 52.06 48.07
-+ GPTQ 4 52.07 47.08 49.83 46.09
-+ AWQ 4 53.58 48.12 50.77 47.43
-+ SPQR 4.63 54.05 49.26 51.99 47.35
-+ GWQ-O (Ours) 4.63 54.06 49.50 52.03 48.07
-+ GWQ-R (Ours) 3.63 53.87 49.32 52.01 49.98
-Table 6: Performance of the Qwen-VL family on the zero-shot tasks on target detection datasets.
-In the three RefCOCO datasets, GWQ is more accurate compared to SPQR for the same bits, achieving less quantization loss.
-QWEN-VL QWEN-VL-chat
-Acc.(%$\uparrow$ Avg_bit test-A test-B val test-A test-B val
-Original 16 92.26 85.34 89.36 92.27 84.51 88.55
-+ GPTQ 4 90.61 83.14 87.13 90.33 82.27 86.67
-+ AWQ 4 91.33 84.51 88.01 91.14 83.61 87.77
-+ SPQR 4.63 92.17 84.87 88.97 91.76 84.20 88.48
-+ SPQR 3.98 91.33 84.51 88.01 91.15 83.61 87.77
-+ GWQ-O (Ours) 4.63 92.18 84.88 88.98 91.78 84.22 88.49
-+ GWQ-R (Ours) 3.98 91.34 84.52 88.03 91.16 83.62 87.79
-Table 7: Comparison of speedup and memory consumption of different quantization methods. GWQ significantly improves the inference speed and reduces the memory required compared to the original model.
-Llama2-7B-hf Llama2-13B-hf
-Methods Avg_bit Speedup($\uparrow$) Memory($\downarrow$) Speedup($\uparrow$) Memory($\downarrow$)
-Original 16 $\times$1 12.8 $\times$1 23.6
-+ GPTQ 4 $\times$2.0 3.1 $\times$2.0 6.7
-+ AWQ 4 $\times$2.1 3.1 $\times$2.1 6.7
-+ SPQR 4.63 $\times$1.2 N/A $\times$1.2 N/A
-+ GWQ (Ours) 4.63 $\times$1.2 4.1 $\times$1.2 7.1
-Language Model.
+Table 4: Results of Llama family on massive multitask language understanding task.
+GWQ outperforms other quantization methods, especially when the average bit-width of the model is significantly lower.
+Accuracy (%)
+Model Methods
+Average Bit
+Calib. Num
+Average STEM Humanities
+Social Sciences
+Other
+Llama-2-7B-hf Original 16 - 45.96 37.04 43.38 51.84 52.44
++OPTQ 4 1024 33.97 25.34 32.41 38.89 39.22
++QuIP 4 1024 34.58 25.90 33.12 39.02 40.29
++GPTQ 4 1024 36.26 27.95 34.55 40.95 41.59
++AWQ 4 512 38.64 30.13 36.85 43.35 44.24
++SPQR-O 4.63 1024 41.34 33.63 39.38 46.65 47.25
++SPQR-R 3.98 1024 38.66 30.40 36.35 43.57 44.33
++GWQ-O 4.63 1 43.36 34.59 40.94 48.32 50.34
++GWQ-R 3.98 1 39.90 31.89 37.54 45.09 45.09
+Llama-2-13B-hf Original 16 - 55.68 44.27 54.43 63.41 60.76
++OPTQ 4 1024 43.32 30.85 39.43 52.59 50.44
++QuIP 4 1024 53.54 31.59 40.22 53.90 51.44
++GPTQ 4 1024 45.47 33.45 41.49 54.46 52.49
++AWQ 4 512 47.35 35.48 43.46 56.1 54.33
++SPQR-O 4.63 1024 49.85 37.59 46.45 58.44 56.93
++SPQR-R 3.98 1024 47.90 35.94 44.23 56.57 54.87
++GWQ-O 4.63 1 52.35 40.82 49.22 61.07 59.34
++GWQ-R 3.98 1 49.04 36.95 45.95 57.35 55.90
+Llama-2-13B-hf Original 16 - 65.01 55.39 56.57 76.87 71.35
++OPTQ 4 1024 51.90 43.59 44.49 61.59 57.93
++QuIP 4 1024 54.12 45.93 46.29 64.95 59.35
++GPTQ 4 1024 56.00 47.84 48.35 66.49 61.35
++AWQ 4 512 57.84 49.24 50.43 68.44 63.24
++SPQR-O 4.63 1024 60.84 51.94 52.55 71.44 67.43
++SPQR-R 3.98 1024 58.11 49.96 50.54 68.49 63.46
++GWQ-O 4.63 1 63.12 54.08 55.94 75.33 70.36
++GWQ-R 3.98 1 60.78 51.48 52.46 71.43 66.54
+Language Modeling Task.
+We first evaluate the perplexity and accuracy of the quantized model on the WikiText2 Merity et al.
+(2016) and C4 Raffel et al.
+(2020) datasets.
 As shown in Tab.
 2 and Tab.
-3, we first evaluate the perplexity of the quantized model on the WikiText2 Merity et al.
-(2016) and C4 Raffel et al.
-(2020) datasets on the Llama2-7B and 13B models.
-Additionally, we assess the models’ ability to zero-shot on these three datasets and compare the results to several baseline methods.
-GWQ achieves a lower average bit-width than other quantization methods by applying more aggressive compression to the non-outlier weights.
-Compared to other quantization approaches, GWQ yields a lower PPL on all datasets under identical bit compression.
-For the zero-shot task, the Acc.
-of the quantized model of GWQ is significantly higher than that of other quantization methods.
-indicating that GWQ results in less performance degradation.
-Notably, under extreme compression, GWQ sustains comparable performance to 4-bit quantization, even at 3-bit.
-Figure 3: Performance of Different Quantized Qwen-VL. GWQ performs well compared to other quantization methods in the zero-shot grounding detection task.
-Table 8: Performance of GWQ at 4bit quantization for different number of calibration samples.
-The quantization loss of the model essentially does not vary with the number of different calibration samples, allowing for low-loss quantization with fewer samples.
-num_sample=1 num_sample=4 num_sample=8 num_sample=12
-[…]
-Llama2-7B-hf 5.51 7.01 5.50 7.01 5.51 7.00 5.50 6.99
-Llama2-13B-hf 4.91 6.48 4.90 6.49 4.91 6.47 4.91 6.48
-Llama-3-8B-hf 6.40 9.21 6.39 9.20 6.38 9.20 6.40 9.20
-Table 9: Performance of GWQ at 4bit quantization for different sample of calibration samples.
-The model quantization loss is essentially invariant to changes in the quantization sample and is less dependent on sample type.
-sample one sample two sample three sample four
-[…]
-Llama2-7B-hf 5.51 7.01 4.91 6.97 5.47 6.97 5.47 6.97
-Llama2-13B-hf 4.91 6.48 4.88 6.47 4.88 6.47 4.88 6.47
-Llama-3-8B-hf 6.40 9.21 5.25 7.75 5.25 7.75 5.25 7.75
-Multimodal Models.
-For the multimodal model, we evaluate the performance of the compressed model in language tasks and multimodal tasks with the Qwen-VL family (Bai et al., 2023a, b).
-Similar to the language models, we evaluate the language generation capabilities of the quantized model on the WikiText2, C4 dataset.
-The multimodal tasks are mainly focused on zero-shot target detection.
+3, at the same average bits, the GWQ compressed model has lower PPL as well as higher Acc compared to the SPQR compressed model.
+Such results show that the quantized model with the same compression rate GWQ performs better in language modeling than the quantized model with SPQR.
+When the average bit-width in the model after GWQ quantization is lower than 4, it still outperforms the 4-bit full-precision quantized model.
+This indicates that GWQ quantization is more robust compared to the full-precision approach.
+For the Qwen-VL family Bai et al.
+(2023a, b) of models whose foundational model is not Llama, the quantized model of GWQ can still maintain strong language modeling ability.
+Thus, it is shown that GWQ can be generalized to different foundational models.
+Grounding Detection.
+The grounding detection tasks are mainly focused on zero-shot target detection.
 We choose the RefCOCO (Kazemzadeh et al., 2014; Yu et al., 2016) dataset to evaluate the zero-shot capability of the quantized model compared to other quantization methods.
@@
 where $correct\_num$ is the number of samples detected correctly and $total\_num$ is the total number of samples.
-As it shown in Tab.
-4 and Tab.
-5, quantized qwen-vl family model of GWQ outperforms the current stat-of-the-arts in both the language task and the zero-shot test, showing better performance.
-The GWQ quantized MLLM performs better on the verbal task compared to other quantization methods, so we will use the quantized model to test it on the visual grounding task.
-As it shown in Tab.
-6, in the zero-shot grounding detection task, both GWQ-quantified Qwen-VL family models with currently outperform the current stat-of-the-arts on most of the datasets, achieving a generalization of the multimodal model, The performance is shown in Fig.
+As shown in Tab.
+3, Compared to SPQR, GWQ quantized same-precision Qwen-VL family models show a significant increase in detection accuracy.
+This indicates that the GWQ method is better suited for grounding detection scenarios compared to SPQR.
+Compared to the full-precision quantization method, the detection ability of the GWQ quantized model is still better than the models quantized by other methods when the average bits of the GWQ quantized model are less than 4.
+This suggests that the GWQ quantized model not only performs well on language modeling tasks but also has the potential to generalize to multimodal scenarios.
+The performance is shown in Fig.
 3, for the task of visual grounding detection of long context, the GWQ post-quantization model is more advantageous than SPQR post-quantization.
-Model Efficiency Comparison.
+Table 5: Evolution of vison-language question answering task datasets.
+GWQ outperforms other quantization methods on both LLaVA and Qwen-VL families.
+Results on ScienceQA dataset, Question classes: NAT is Natural Science, SOC is Social Science, LAN is Language Science, TXT is Text Context, IMG is Image context, NO = No context.
+Accuracy (%)
+Model Methods
+Average Bit
+Calib. Num
+Average NAT SOC LAN TXT IMG NO
+LlaVA-V1.5 Original 16 - 62.94 60.98 63.10 67.27 62.99 63.24 60.04
++OPTQ 4 1024 49.91 48.01 50.12 55.84 48.96 48.92 47.63
++QuIP 4 1024 51.76 49.98 51.92 57.78 50.69 50.63 49.58
++GPTQ 4 1024 52.69 51.01 52.95 58.64 51.73 51.71 50.15
++AWQ 4 512 54.66 53.97 55.37 61.36 54.70 54.66 52.29
++SPQR-O 4.63 1024 58.08 56.09 58.03 62.84 57.88 57.93 55.68
++SPQR-R 3.98 1024 55.23 53.99 55.36 60.36 54.68 54.67 52.31
++GWQ-O 4.63 1 60.55 58.89 60.68 64.98 60.24 60.56 57.94
++GWQ-R 3.98 1 56.99 55.01 57.10 61.82 56.78 56.84 54.36
+Qwen-VL-chat Original 16 - 68.89 73.20 65.12 68.77 71.07 66.84 68.35
++OPTQ 4 1024 55.99 55.14 58.46 51.72 54.98 60.76 54.85
++QuIP 4 1024 57.75 56.98 60.02 53.58 56.82 62.34 56.76
++GPTQ 4 1024 58.65 57.99 60.64 54.25 57.84 63.24 57.96
++AWQ 4 512 61.39 60.93 63.32 57.15 60.55 65.90 60.46
++SPQR-O 4.63 1024 64.03 67.56 59.91 63.24 66.25 64.02 63.21
++SPQR-R 3.98 1024 61.38 60.92 63.33 57.14 60.56 65.91 60.43
++GWQ-O 4.63 1 67.39 70.16 62.20 65.91 68.72 66.14 65.85
++GWQ-R 3.98 1 64.03 67.66 59.98 62.12 68.33 63.96 62.14
+Qwen-VL Original 16 - 66.70 63.55 66.48 69.02 67.91 66.25 66.99
++OPTQ 4 1024 54.03 50.89 52.97 57.03 56.69 52.98 53.64
++QuIP 4 1024 55.80 52.76 54.81 58.65 58.23 54.94 55.38
++GPTQ 4 1024 56.64 53.57 55.68 59.49 59.07 55.86 56.14
++AWQ 4 512 59.17 56.02 58.32 61.90 61.01 58.72 59.02
++SPQR-O 4.63 1024 61.81 58.64 61.93 64.14 62.43 61.76 61.98
++SPQR-R 3.98 1024 59.17 56.01 58.34 61.92 60.99 58.71 59.04
++GWQ-O 4.63 1 63.87 60.43 63.59 66.34 64.94 63.46 64.46
++GWQ-R 3.98 1 61.37 57.53 60.84 63.74 63.01 61.44 61.87
+Table 6: efficiency comparison between different quantization methods. GWQ is not the best compared with other quantization methods.
+Average Calibration Llama-2-7B-hf Llama-2-13B-hf Llama-3-8B-hf
+bit number Speedup ($\uparrow$) Memory ($\downarrow$) Speedup ($\uparrow$) Memory ($\downarrow$) Speedup ($\uparrow$) Memory ($\downarrow$)
+Original 16 - $\times$1.00 12.83G $\times$1.00 23.63G $\times$1.00 14.57G
++OPTQ 4 1024 $\times$1.90 3.21G $\times$1.91 6.9G $\times$1.90 3.84G
++QuIP 4 1024 $\times$1.98 3.11G $\times$1.99 6.71G $\times$1.98 3.61G
++GPTQ 4 1024 $\times$2.00 3.11G $\times$2.01 6.71G $\times$2.00 3.61G
++AWQ 4 512 $\times$2.01 3.10G $\times$2.00 6.72G $\times$2.01 3.61G
++SPQR-O 4.63 1024 $\times$1.28 N/A $\times$1.27 N/A $\times$1.28 N/A
++SPQR-R 3.98 1024 $\times$1.26 N/A $\times$1.25 N/A $\times$1.26 N/A
++GWQ-O 4.63 1 $\times$1.23 4.16G $\times$1.22 7.13G $\times$1.23 4.84G
++GWQ-R 3.98 1 $\times$1.22 4.16G $\times$1.21 7.13G $\times$1.22 4.84G
+Massive Multitask Language Understanding.
+In this section, we will discuss the ability of Llama-2 and Llama-3 families using different quantization methods in massive multitask language understanding with MMLU dataset Hendrycks et al.
+(2021a, b).
+As shown in Tab.
+5, compared with SPQR, GWQ quantized Llama-2 and Llama-3 families show a substantial improvement in multimodal comprehension accuracy.
+This indicates that GWQ is more suitable for application in complex scenarios than SPQR.
+Compared with the full-precision quantization method, when the average bits of the GWQ quantized model are less than 4, the GWQ quantized model still outperforms the models quantized by the other methods in language understanding ability.
+The results show that our method maintains proficiency in single-task language modeling and also exhibits strong adaptability in complex multi-task scenarios.
+Table 7: Effect of the number of calibration set to the quantization.
+Different numbers of calibration set affect a little to quantifies models’ perform.
+All calibration set are from Red Pajama1 dataset.
+sample_number=1 sample_number=4 sample_number=16 sample_number=64 sample_number=512 sample_number=1024
+[…]
+Llama2-7B-hf 5.48 6.99 5.49 6.98 5.47 6.98 5.45 6.96 5.46 6.97 5.46 6.96
+Llama2-13B-hf 4.90 6.48 4.91 6.47 4.90 6.47 4.89 6.46 4.89 6.48 4.89 6.48
+Llama3-8B-hf 6.41 9.15 6.43 9.17 6.42 9.16 6.41 9.15 6.40 9.14 6.40 9.14
+Qwen-VL 8.24 10.10 8.23 10.10 8.25 10.11 8.23 10.09 8.23 10.09 8.23 10.09
+Qwen-VL-chat 9.86 11.88 9.86 11.88 9.85 11.86 9.85 11.87 9.85 11.87 9.85 11.86
+Table 8: Effect of different calibration set to quantization results. different calibration set effects a little to quantization results, but special calibration set of task affects a lot.
+WikiText2 C4 Red Pajama1 Red Pajama2 AWQ’s Calibration
+[…]
+Llama2-7B-hf 5.46 6.96 5.47 6.92 5.48 6.99 5.48 6.98 5.80 7.03
+Llama2-13B-hf 4.85 6.47 4.88 6.85 4.90 6.48 4.91 6.47 5.99 6.51
+Llama3-8B-hf 6.38 9.14 6.40 9.10 6.41 9.15 6.41 9.16 6.49 9.56
+Qwen-VL 8.22 10.08 8.23 10.05 8.24 10.10 8.23 10.11 8.31 10.17
+Qwen-VL-chat 9.84 11.89 9.86 11.86 9.86 11.88 9.87 11.87 9.93 11.94
+Vision-Language Question Answering.
+The vision-language question answering (VQA) tasks aim to assess the model’s ability to generate accurate responses to questions based on visual and textual inputs.
+We evaluate the performance of GWQ and other quantitative methods applied to LlaVA-V1.5 Liu et al.
+(2024), Qwen-VL Bai et al.
+(2023a), and Qwen-VL-chat Bai et al.
+(2023a) models on the ScienceQA dataset Lu et al.
+(2022).
+As shown in Table 4, similar to the previous task, the performance of the GWQ quantized model achieved SOTA on VQA accuracy, compared to the other methods.
+This suggests that GWQ is highly effective for vision-language question-answering tasks.
+Even when the average bits of the GWQ quantized models are below 4, the GWQ quantized models still outperform models quantized by other methods.
+This indicates that the GWQ quantized models not only maintain strong performance in language modeling but also generalize well to complex multimodal tasks like VQA.
+Model Efficiency Comparison. In this section, we will discuss the impact of different quantization methods on model acceleration and compression effects.
 For the model throughput (tokens/s), the speedup rate can be obtained by calculating the time required to generate 512 tokens after quantization and comparing it with the original model.
-For the model inference overhead (G), we use the peak value when the model generates 512 tokens as the maximum inference memory by recording it.
-As illustrated in Tab.
-7, the GWQ quantized model can achieve about 1.2$\times$ speedup relative to the original model, which is on par with the current mixed-accuracy method SPQR.
-Compared to the full-precision quantization method, GWQ results in higher inference latency due to floating-point conversion during inference.
-Compared to the original model, GWQ consumes less memory during inference, which makes it easier to deploy on edge devices.
-Figure 4: Effect of calibration sample on quantification.
-For different numbers of calibration sets, the PPL of the quantized model tends to be essentially constant on WikiText2 and C4.
-Therefore, the effect of the number of calibration set on the quantization effect is negligible.
+For the model inference overhead (G), we use the peak memory when the model generates 512 tokens as the maximum inference memory by recording it.
+As shown in Tab.
+6, the model after quantization using GWQ 4bit achieves about 1.2$\times$ of speedup compared to the original model.
+However, compared to SPQR, the acceleration effect of GWQ is not significant.
+We believe that the insignificant effect of GWQ is due to the fact that the outlier values saved by GWQ search are sparse, which is not favorable for inference calculation.
+Compared with the full-precision quantization methods, GWQ and SPQR are not advantageous due to their hardware adaptability.
 4.3 Ablation Study
-In this section, we examine the impact of varying the number of calibration samples on the performance of our method.
-Tab.
-8 presents the results for the calibration sample sizes 1, 4, 8 and 12, evaluated on the WikiText2 and C4 datasets.
-Tab.
-8 and Fig.
-4 indicate that increasing the number of calibration samples has minimal impact on the PPL of the quantized model.
-This suggests that the size of the calibration set has little influence on the quantization performance of GWQ.
-Compared to most of the current quantization methods, which require large datasets, GWQ has a lower calibration cost and can achieve calibration with fewer samples.
-To further investigate the impact of different calibration samples, we selected four distinct samples, each of which was used as a calibration set to quantize the model.
-According to Tab.
-9, when different calibration samples are used to quantize the same model, the resulting PPL values on the same dataset remain similar, indicating that the choice of calibration samples has a negligible effect on the quantization results.
-Unlike AWQ, GWQ can be well adapted to the calibration sets of different samples, and the quantized model performs well even if the calibration sets differ greatly.
+In this section, we test the number of checksets and the effect of different checksets on our quantization method and justify our choice of parameters.
+Effect on Number of Calibration.
+We select different number of samples from the Red Pajama dataset as calibration samples respectively, by taking the gradient mean and searching for the corresponding outlier of the model.
+We tested the performance of the model after quantizing different numbers of samples on the WikiText2 and C4 datasets.
+As shown in Tab.
+7, we find that the model effect fluctuates slightly when the number of calibration samples is small, but it does not affect the final quantitative results very much.
+At the same time, when the number of calibration samples is increased, we find that the effect of the quantized model is similar to that when there is only one sample.
+Therefore, from the consideration of data efficiency and method robustness, we believe that it is optimal to choose a sample of 1.
+Effect on Different Calibration.
+We choose different calibration data Lin et al.
+(2024a); Weber et al.
+(2024); Merity et al.
+(2016); Raffel et al.
+(2020) to calibrate the Llama families model Dubey et al.
+(2024); Touvron et al.
+(2023b) and test the PPL of the quantized model on the WikiText Merity et al.
+(2016) and C4 Raffel et al.
+(2020) datasets.
+As shown in Tab.
+8, different types of calibration samples do not have a significant effect on the quantized model.
+We guess that the impact of the calibration data on the quantization is less due to the fact that when collecting the gradients, we used the gradient mean as the outlier indexing basis.
+However, for specialized calibration sets, such as using the C4 calibration set for quantization calibration, and when testing on the C4 calibration set, the model will show better results.
+At the same time, we believe that the quantization effect is related to the complexity of the checksum set.
+When WikiText is used as the check set, its quantized model effect has a slight decrease compared to the rest of the models.
+This may be due to the fact that the language structure of the WikiText check set is relatively simple, which cannot realize more accurate calibration.
 5 Conclusion
-This paper introduces GWQ, a weight-only post-training mixed-precision quantization approach based on first-order gradients.
-GWQ preserves the weights corresponding to the top 1% of outliers at FP16 precision, while storing the remaining non-outlier weights in a lower-bit format.
-GWQ found that it is easier to obtain more accurate outliers using the first-order gradient to locate the sensitive weights of the pre-trained model than the current search for model outliers via the Hessian matrix.
-GWQ achieves the lowest quantization loss compared to current state-of-the-arts methods by utilizing only a single calibration sample.
-Considerable results are also obtained by GWQ on multimodal models, which can achieve better results on most of the zero-shot target detection tasks compared to the current stat-of-the-art methods.
-GWQ achieves 1.2$\times$ inference acceleration compared to the original model internship and uses less memory in the inference process.
+This paper introduces GWQ, a weight-only post-training mixed-precision quantization based on first-order gradients.
+GWQ preserves the weights corresponding to the top 1% of outliers at FP16 bit-width while storing the remaining non-outlier weights in lower-bit.
+It found that it is easier to obtain more accurate outliers using the first-order gradient than Hessian matrix.
+GWQ achieves the lowest quantization loss compared to current SOTA methods with only one single calibration sample.
+In both text-only tasks and multimodel tasks, LLMs quanted by GWQ perform better than other quantization methods, which indecate that GWQ could apply to different models and tasks.
+Also, GWQ achieves 1.2$\times$ inference acceleration compared to the original model internship and uses less memory in the inference process.
 6 Limitation
@@
 However, in models such as OPT, the activation function is ReLU, and the correct gradients cannot be computed because of gradient vanishing.
-Additionally, backpropagation demands substantial memory resources, a single A100 80G GPU is insufficient to handle the large computational requirements.
+Additionally, backpropagation demands substantial memory resources, and a single A100 80G GPU is insufficient to handle the large computational requirements.
 Furthermore, since GWQ employs mixed-precision quantization, it is less hardware-friendly compared to methods like AWQ, resulting in higher inference latency after quantization.
 We will continue to refine and optimize the GWQ methodology.
+Appendix A Appendix
+A.1 Next Token Prediction Accuracy
+The accuracy evaluation of for WikiText2 and C4 datasets is based on a causal language modeling task, where given a ground truth sequence of tokens $\mathbf{y}=\{\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{T}\}$ and the predicted tokens $\hat{\mathbf{y}}=\{\hat{\mathbf{y}}_{1},\hat{\mathbf{y}}_{2},\ldots,\hat{\mathbf{y}}_{T}\}$, the accuracy is computed by shifted ground truth $\vec{\mathbf{y}}$
+$\vec{\mathbf{y}}=\mathbf{y}[:,1:]$ (15)
+and the most probable token of the model output logits $\mathbf{L}$, where $\mathbf{L}$ is shifted to remove the last token’s predictions
+$\mathbf{L}_{\text{shifted}}=\mathbf{L}[:,:-1,:]$ (16)
+and predictions $\hat{y}$ are
+$\hat{\mathbf{y}}_{t}=\arg\max_{v}\mathbf{L}_{t,v}$ (17)
+where $v$ is the vocabulary size. Finally, accuracy is
+$\mathrm{Acc}=\frac{N_{\hat{y_{t}}=\overrightarrow{y_{t}}}}{T},$ (18)